NvidiaIntely (nvidiaintely):

In a nuclear reaction, the energy released is equal to 8.1 x 1016 joules. Calculate the mass lost in this reaction. (1 J = 1 kg m2/s2) 7.2 x 1032 kg 2.7 x 108 kg 9.0 x 10-1 kg 7.2 x 10-16 kg

OpenStudy (da01159):

What is the Question?

NvidiaIntely (nvidiaintely):

Chemistry.

OpenStudy (da01159):

so 7.2 is what unit

NvidiaIntely (nvidiaintely):

kg

OpenStudy (kevin):

Have you calculate it @NvidiaIntely

OpenStudy (kevin):

I got E = m x c^2 8.1 * 10^16 = m * (3*10^8)^2 m = 1.11 Idk if I calculate it wrong

OpenStudy (kevin):

Maybe we can go to another question while waiting for other response. lol

NvidiaIntely (nvidiaintely):

hmmm.. Well let me check Brainly. Here: Use the formula E=mc^2 energy given=8.1 x 10^16 joules speed of light = 3.00 × 10^8 m/s plug in the values we'll get mass=9.0 x 10-1 kg Again no explaining...

OpenStudy (kevin):

Ok, than it will be 9.0 x 10-1 kg

NvidiaIntely (nvidiaintely):

Yeah, okay then next question...